# Character-Adapter: Prompt-Guided Region Control for High-Fidelity Character Customization

### v2

@@ 3 Methods
3.1 Preliminaries
Latent diffusion model. Latent diffusion models (LDMs) [28] consist of three components: a text encoder, a variational autoencoder (VAE) [25], and a U-Net. The U-Net aims to predict noise during the diffusion process, as follows:
$L_{LDM}:=\mathbb{E}_{\varepsilon(x),\epsilon\sim\mathcal{N}(0,1),t}[\parallel\epsilon-\epsilon_{\theta}(z_{t},t)\parallel]_{2}^{2}]$ (1)
where $z_{t}$ is obtained from the encoder $\mathcal{E}$. Cross-attention in LDMs. Cross-attention mechanisms are used to augment the underlying U-Net backbone, thereby transforming DMs into more flexible conditional image generators, which are effective for learning attention-based models for a variety of input modalities [29]. In the original Stable Diffusion, text prompts are encoded by CLIP [30] to obtain text features, which are fed as conditions into the U-Net backbone of the diffusion model. Specifically, the output after the cross-attention mechanism in the Stable Diffusion are noted as:
where $Q=zW_{q}$,$K=c_{t}W_{k}$,$V=c_{t}W_{v}$, $z$ denotes latent vectors obtained from the encoder $\mathcal{E}$, $c_{t}$ denotes text features obtained from text encoder (CLIP in Stable Diffusion).
Image Prompt Adapter. IP-Adapter proposes a decoupled cross-attention strategy to support conditional image generation by introducing an image cross-attention mechanism [9] analogous to the original cross-attention module in Stable Diffusion [28]. This mechanism seamlessly integrates image prompts with text prompts to guide the text-to-image generation process. Owing to the decoupled nature of the image and text cross-attention mechanisms, the proposed approach can be readily integrated into existing Stable Diffusion models without additional training. The decoupled cross-attention can be formulated as:
@@ 3.1 Preliminaries
where $K^{\prime}=c_{i}{W}^{\prime}_{k}$,$V^{\prime}=c_{i}{W}^{\prime}_{v}$, $c_{i}$ represents the image features extracted from the CLIP image encoder, and $Z^{new}$ denotes the new latent representation obtained by conditioning on both image and text inputs. Notably, ${W}^{\prime}_{k}$ and ${W}^{\prime}_{v}$ are the only trainable parameters in IP-Adapter.
@@ 3.2 Framework of Character-Adapter
In this section, we introduce Character-Adapter, a novel framework designed for high-fidelity consistent character generation. As illustrated in Fig. 2, we first segment a reference character into several parts using prompt-guided segmentation. Subsequently, We obtain the attention maps of the target image layout using the same approach. Finally, we utilize dynamic region-level adapters to achieve detailed and consistent character generation.
@@ 3.2.1 Prompt-guided segmentation
We hypothesize that directly passing the entire reference image into the image encoder leads to inadequate extraction of detailed image features pertaining to the given character [9], as evidenced by the examples shown in Fig. 4. To address this problem, we propose a prompt-guided segmentation module that decomposes the given character into separate regions, namely the face, upper body, and lower body. The ablation study presented in Sec. 4.4 proves our assumption and substantiates the efficacy of the proposed prompt-guided segmentation in facilitating comprehensive image feature extraction.
@@
When given a user-input prompt $P$, for instance, “a boy standing in a library, wearing green jacket and blue pants”, we complete the prompt with detailed region descriptions:
$P_{C}=[P_{G};P_{F};P_{U};P_{L}]=W:=[w_{1};...;w_{g};w_{g+1}...w_{f}...w_{l}],\quad l>u>f>g,$ (4)
@@
where $P_{G}$ denotes the user-given prompt, $P_{F}$ indicates the prompt for the face region (“a boy”), $P_{U}$ specifies the prompt for the upper body (“green jacket”), and $P_{L}$ relates to the prompt for the lower body (“blue pants”). $W:$ represents the sequence of words in the constructed prompt $P_{C}$.
Subsequently, We use $P_{C}$ to generate a layout image with Stable Diffusion, from which we derive the attention map corresponding to each word and the latent representation of U-Net. To be specific, let us consider a 2D layout image latent, denoted as $z_{t}$, at timestep $t$. The attention map for the $i$-th word $w_{i}$, in relation to the image latent at the $k$-th U-Net block, can be represented as follows:
$A_{i,t}^{(k)}=F(w_{i},z_{t}^{(k)})=softmax((W_{q}^{(k)}\cdot z_{t}^{(k)})(W_{k}^{(k)}\cdot TE(w_{i}))),$ (5)
@@
where $A$ denotes the attention map, $W_{q}$ and $W_{k}$ are projection matrices in the attention module, and $TE$ corresponds to a language model that encodes text prompts into text embeddings.
To obtain precise attention maps and masks, we integrate the attention maps of all $K$ layers to a uniform dimension through interpolation. Subsequently, we calculate the similarity $S_{i,t}$ as follows:
$S_{i,t}=\sum_{k=1}^{K}Interplot(A_{i,t}^{(k)},w,h),$ (6)
where $w$ and $h$ represent the width and height of the noised image’s latent, respectively.
@@
Since each semantic region prompt comprises multiple words, such as “green jacket”, we aggregate each attention map of these words, denoted by $S_{i,t}$, in order to establish the correlation between the region prompt and the corresponding semantic image region:
$S_{r,t}=\mathop{\max}\limits_{r_{begin}<i<r_{end}}\sum_{k\in[1,K]}Interplot(A_{i,t}^{(k)},w,h),$ (7)
@@
where $r_{begin}$ and $r_{end}$ represent the word indices corresponding to the beginning and end of the region prompt, respectively.
Similarly, given a reference character, we follow Eq.  5 and utilize attention maps to extract specific adapter condition regions. First, we add noise to the latent representation of the reference image $z_{0}^{(R)}$ in $t$ steps as the following:
$z_{t}^{(R)}=\sqrt{\overline{\alpha}_{t}}z_{0}^{(R)}+\sqrt{1-\overline{\alpha}t}\epsilon_{t}.$ (8)
Following Eq.  5, the corresponding attention map $S_{r,t}^{R}$ between the reference image $R$ and the prompt $P_{C}$ is obtained. We then use attention maps between the prompt and reference image to achieve a more precise and fine-grained segmentation of the reference image:
$y_{begin}^{r}=\mathop{\arg\min}\limits_{y}(S_{r}^{R}(x,y)>\gamma_{2}),\quad x_{begin}^{r}=0,$ (9)
$y_{end}^{r}=\mathop{\arg\max}\limits_{y}(S_{r}^{R}(x,y)>\gamma_{2}),\quad x_{end}^{r}=width,$ (10)
we ultimately employ $CR_{r}(x,y)$, defined within the ranges $x\in[x_{begin}^{r},x_{end}^{r}]$ and $y\in[y_{begin}^{r},y_{end}^{r}]$, as the semantic region of the reference image, segmented through prompt guidance.
@@ 3.2.2 Dynamic region-level adapters
Another limitation of existing consistent character generation methods is concept fusion, which we posit stems from the text encoder’s inability to adequately encapsulate compositional concepts. Employing full tokens for multi-concept generation may lead to concept confusion [15]. This diminishes the guidance of these features toward the corresponding regions of the latent vectors, as substantiated in Sec. 4.4. To mitigate this limitation, we propose a dynamic region-level adapters module to alleviate concept fusion during the inference process.
Region-level adapters. Initially, we employ mask-based multi-adapters to integrate the semantic guidance from different regions. The mask at resolution $(x,y)$ of each region can be obtained as following:
@@
$M_{r}(x,y)=Binary(S_{r}(x,y))=\left\{\begin{array}[]{l}1\quad\quad S_{r}(x,y)>\gamma_{1},\\ 0\quad\quad S_{r}(x,y)<=\gamma_{1}.\\ \end{array}x\in[1,w],y\in[1,h]\right.,$ (11)
where $S_{r}(x,y)$ is the attention map of specific region of the layout image. $\gamma_{1}$ is the threshold.
As these adapters represent different semantic regions of a reference image, for example, the face region is used to guide the character’s appearance and hairstyle. A natural approach is to fuse the different semantics using mask guidance, the masked attention map now reads:
$A_{t}(x,y)=\sum_{r}(A_{r,t}^{(CR)(x,y)}\cdot M_{r}(x,y))+\prod\limits_{r}(1-M_{r}(x,y))\cdot A_{P,t}(x,y),$ (12)
where $A_{r,t}^{(CR)}$ refers to the attention map between the reference image region $CR_{r}$ and the prompt $P_{C}$ using Eq. 7, $M_{r}(x,y)$ refers to the masks obtained from the layout image as shown in Eq. 11, $A_{P,t}(x,y)$ is the attention map derived solely from the prompt condition.
@@
Attention dynamic fusion. Mask-based regional-level adapters face the challenge of obtaining region masks that facilitate generation adhering to the layout while remaining faithful to the reference image. We observe that an inference time of $T=650$ yields masks whose shapes strike a balance between the reference and layout masks (Fig. 3). However, the coarse nature of these hard-label masks adversely impacts the generation of well-defined boundaries, as evidenced in Sec. 4.4. To address this, we propose an attention dynamic fusion module, characterized as a soft-label mask, to integrate the region-level adapters and produce coherent image generation through $A_{t}(x,y)$, defined as follows:
$A_{t}(x,y)=\sum_{r}(A_{r,t}^{(CR)(x,y)}\cdot softmax(S_{r}(x,y))+softmax(1-\sum_{r}S_{r}(x,y))\cdot A_{P,t}(x,y).$ (13)
3.2.3 Multi-character consistency
@@
In addition to achieving remarkable performance in terms of single-character consistency, Character-Adapter also demonstrates remarkable performance for multi-character generation. Similar to single-character generation, we utilize prompt-guided segmentation to automatically localize multiple reference characters. The prompt includes a global prompt and regional prompts for $N$ character is noted as:
$P_{C}=[P_{G};P_{F}{}_{1};P_{U}{}_{1};P_{L}{}_{1};...P_{F}{}_{N};P_{U}{}_{N};P_{L}{}_{N}].$ (14)
@@
0:  Prompt $P_{C}$, $N$ reference character images $R_{i},i\in[1,N]$, a diffusion network $D_{\theta}$
1:  use prompt $P_{C}$ to obtain attention map $A_{i,t}^{(k)}$ via Eq. 5 and correlation map $S_{r,t}$ via Eq. 7
2:  for i in [1, N] do
3:     add noise and perform diffusion process to get correlation map $S_{r,t}^{R^{(i)}}$ via Eq. 5 Eq. 7
4:     get region adapter inputs $CR_{r}^{(i)}$ via Eq. 9 and Eq. 10
5:  end for
6:  fuse the adapters to get attention map $A_{t}$ via Eq. 13
6:  target image $T_{I}=D_{\theta}(A_{t})$
@@ 4.1 Experimental setup
Implementation details. Character-Adapter is a plug-and-play framework and can be transplanted to any diffusion backbone. Hence, we conduct all experiments on SD v1.5 [28] with classifier-free guidance as 7 and 20-timestep Euler A sampling. The mask threshold $\gamma_{1}$ and $\gamma_{2}$ are all set as 0.8. More details are provided in Sec. A.1
@@ 4.2 Quantitative comparison
We conduct experiments on Character-Adapter and previous SOTA works, quantitative results are illustrated in Table 1. state-of-the-art performance in zero-shot character consistency generation, with an improvement of 24.8% on the CLIP-I and DINO scores compared to previous state-of-the-art methods in both single and multi character generation. In comparison with fine-tuning-based approaches, Character-Adapter can achieve comparable character consistency while offering over 70 times improvement in computational efficiency, as it requires no additional training. Moreover, Character-Adapter achieves the best performance in terms of text-image alignment with an improvement of 3.5%.
Table 1: Quantitative results (%) of Character-Adapter with other methods. Evaluations are conducted for both single and multiple character generation. The best results are highlighted in bold.
@@ 4.3 Qualitative comparison
As shown in the multi-character generation example (Fig. 4 6th row), methods like T2I-Adapter, InstantID, Reference-only, and FastComposer fail to preserve the reference characters’ attire. IP-Adapter encounters concept fusion and detail omission issues, attributable to using an image-level adapter, leading to inadequate feature extraction and representation. While fine-tuning-based models like LoRA align with text prompts, they lack details of reference characters and cause concept fusion in multi-character generation. In contrast, Character-Adapter generates high-fidelity multi-character images with intricate details while preserving text-image alignment. Overall, the qualitative results substantiate Character-Adapter’s significant improvements in high-fidelity character generation. Additional results are provided in Sec. A.5.

### v4

@@ 3 Methods
3.1 Preliminaries
Cross-attention in LDMs. Cross-attention mechanisms are used to augment the underlying U-Net backbone, thereby transforming DMs into more flexible conditional image generators, which are effective for learning attention-based models for a variety of input modalities [28]. Specifically, the output after the cross-attention mechanism in the Stable Diffusion are noted as:
where $Q=zW_{q}$,$K=c_{t}W_{k}$,$V=c_{t}W_{v}$, $z$ denotes latent vectors obtained from the encoder $\mathcal{E}$, $c_{t}$ denotes text features obtained from text encoder.
@@ 3.1 Preliminaries
Image Prompt Adapter. IP-Adapter proposes a decoupled cross-attention strategy to support conditional image generation by introducing an image cross-attention mechanism [9] analogous to the original cross-attention module in Stable Diffusion. This mechanism seamlessly integrates image prompts with text prompts to guide the text-to-image generation process. The decoupled cross-attention can be formulated as:
@@
where $K^{\prime}=c_{i}{W}^{\prime}_{k}$,$V^{\prime}=c_{i}{W}^{\prime}_{v}$, $c_{i}$ represents the image features extracted from the CLIP image encoder, and $Z^{new}$ denotes the new latent representation obtained by conditioning on both image and text inputs.
@@ 3.2 Framework of Character-Adapter
In this section, we introduce Character-Adapter, a novel framework designed for high-fidelity consistent character generation. We introduce prompt-guided segmentation to ensure fine-grained regional features of reference characters and dynamic region-level adapters to mitigate concept confusion. As illustrated in Fig. 2, we first segment a reference character into several parts using prompt-guided segmentation. Subsequently, We obtain the attention maps of the target image layout using the same approach. Finally, we utilize dynamic region-level adapters to achieve detailed and consistent character generation.
@@ 3.2.1 Prompt-guided Segmentation
Directly passing the entire reference image into the image encoder can lead to inadequate extraction of detailed image features related to the given character. It potentially causes identity degradation and compromises detailed preservation. To address this, one approach is to divide the entire reference image into essential parts, enabling the image encoder to focus on specific regions of an image for comprehensive feature extraction. We propose a prompt-guided segmentation module, localizing essential image regions based on text prompts, thereby facilitating adequate image feature extraction.
@@
Considering the fundamental aspects of a character, we categorize it into three parts: the face, upper body, and lower body. When given a user-input prompt $P$, for instance, “a boy standing in a library, wearing green jacket and blue pants”, we complete the prompt with detailed region descriptions:
$P=\{P_{m}\}=\{w_{i}\},m=F,U,L\quad i=1,...,N,$ (3)
@@
where $N$ refers to the number of words in the prompts, $P_{F}$ indicates the prompt for the face region (“a boy”), $P_{U}$ specifies the prompt for the upper body (“green jacket”), and $P_{L}$ relates to the prompt for the lower body (“blue pants”). $W:$ represents the sequence of words in the constructed prompt $P$.
Previous works prove the cross-attention maps in the diffusion model contains semantic information of a latent image. Character-Adapter starts with this well-explored finding, seeking to use cross-attention maps between specific text prompts and the latent image to localize the three part of a character. Specifically, given a prompt $P$, the attention map between each word $w_{i}$ and the image latent of $k$-th layer of the UNet can be defined as the following:
$\begin{split}A_{i,t}^{(k)}&=F(w_{i},z_{t}^{k})\\ &=softmax((W_{q}^{(k)}\cdot z_{t}^{k})(W_{k}^{k}\cdot\mathcal{E}(w_{i}))),\end{split}$ (4)
@@
where $A_{i,t}^{(k)}$ denotes the attention map between the word $w_{i}$ and the image latent $z_{t}$ in the $k$-th layer of UNet, $W_{q}$ and $W_{k}$ are projection matrices, and $\mathcal{E}$ corresponds to a language model that encodes text prompts into text embeddings. To obtain precise attention maps, we integrate the attention maps of all $K$ layers in the UNet through interpolation to obtain the final attention map $A_{i,t}$ between $w_{i}$ and the image latent $z_{t}$ as the following:
$A_{i,t}=\sum_{k=1}^{K}A_{i,t}^{(k)},$ (5)
where $K$ denotes the total number of layers in the UNet.
@@
A region prompt, denoted as $P_{m}$ in Eq. 3 contains several words (“green jacket”). Consequently, we aggregate the attention maps associated with these words, denoted by $A_{i,t}$ in Eq. 5, to obtain the cross-attention map between each region prompt and the latent image:
$A_{r,t}=\mathop{\max}\limits_{r_{begin}<i<r_{end}}A_{i,t},$ (6)
@@
where $r_{begin}$ and $r_{end}$ represent the word indices corresponding to the beginning and end of the region prompt, respectively. Therefore, we utilize the attention maps $A_{r,t}$ to partition an image into each region part $I_{P}$ mentioned in Eq. 3:
$y_{1}=\mathop{\arg\min}\limits_{y}(A_{r}(x,y)>\gamma_{1}),\quad x_{begin}^{r}=x_{1},$ (7)
$y_{2}=\mathop{\arg\max}\limits_{y}(A_{r}(x,y)>\gamma_{1}),\quad x_{end}^{r}=x_{2},$ (8)
$I_{P}=[x_{1},x_{2},y_{1},y_{2}]$ (9)
where $x$,$y$ represents the width and the height of the original image, $\gamma_{1}$ is set to 0.8 in our paper.
As illustrated in Stage 1 in Fig. 2, a reference image $I_{R}$ is segmented into three main parts by using Prompt-guided Segmentation. Specifically, we add noise to the latent representation $z_{0}^{R}$ of the reference image $I_{R}$ in $t$ steps as the following:
$z_{t}^{R}=\sqrt{\overline{\alpha}_{t}}z_{0}^{R}+\sqrt{1-\overline{\alpha}t}\epsilon_{t}.$ (10)
Following Eq. 3-9, a reference image $I_{R}$ is ultimately split into three parts: $I_{P_{F}}$,$I_{P_{U}}$,$I_{P_{L}}$.
@@ 3.2.2 Dynamic Region-level Adapters
Given that cross-attention layers are effective in controlling the appearance of the generated image, IP-Adapter proposes an image prompt adapter to inject the feature of a reference image into the diffusion model. However, when multiple image conditions are added to the diffusion model, it becomes unable to generate an image that preserves each of the conditional features, leading to concept fusion. To mitigate this limitation, we propose a dynamic region-level adapters module, using specific regions’ layout to alleviate concept fusion during the inference process.
Upon obtaining each region of the reference image $I_{P_{F}}$,$I_{P_{U}}$,$I_{P_{L}}$ mentioned in Eq. 3, we mainly inject them into diffusion model using a region-level adapters and Attention Dynamic Fusion module.
0:  Prompt $P$, $N$ reference character images $I_{i},i\in[1,N]$, a diffusion network $D_{\theta}$
1:  get the regions of the reference image via Eq. 9
2:  for i in [1, N] do
3:     use prompt $P_{m}$ to obtain attention map $A_{i,t}^{k}$ via Eq. 4 and correlation map $A_{r,t}$ via Eq. 6
4:     get region adapter outputs $I_{P}$ via Eq. 15
5:  end for
6:  fuse the adapters into U-Net to get the hidden state via Eq. 2
6:  target image $T_{I}=D_{\theta}(A_{t})$
Region-level Adapters. Region-level Adapters consist of three individual IP-Adapters, with each IP-Adapter being responsible for injecting features of one region of the reference image. Obtaining the specific layout regions corresponding to each region of the reference image is crucial. In our task, we generate an image with a given prompt and a reference image. During the denoising process, the cross-attention maps between the region prompts and the image latent can be obtained by using Prompt-guided Segmentation mentioned in Eq.3-6. These attention maps can be served as the masks of three regions of the layout image, as illustrated in Stage 2 in Fig. 2. The masks at $(x,y)$ of each region can be obtained as following:
@@
$\begin{split}M_{r,t}(x,y)&=Binary(A_{r,t}(x,y))\\ &=\left\{\begin{array}[]{l}1\quad\quad A_{r,t}(x,y)>\gamma_{2},\\ 0\quad\quad A_{r,t}(x,y)<=\gamma_{2}.\\ \end{array}x\in[1,w],y\in[1,h]\right.,\end{split}$ (11)
where $A_{r}(x,y)$ is the attention map obtained by Eq. 6. $\gamma_{2}$ is the threshold, we set it as 0.8 in our paper.
Since these adapters are tasked with handling distinct regions during image generation, such as using the face adapter to influence the character’s appearance and hairstyle, a logical method involves merging the various adapters through mask guidance. The current masked attention map between a region prompt $P_{m}$ a latent image $z_{t}$ is as follows:
$A_{r,t}^{R}=\sum\limits_{x=0}^{w}\sum\limits_{y=0}^{h}(A_{r,t}(x,y)\cdot M_{r}(x,y)),$ (12)
where $A_{r,t}(x,y)$ can be obtained from Eq. 6 using $P_{m}$. $w$,$h$ refers to the resolution of the latent image $z_{t}$. Subsequently, the current masked attention map is sent into the s elf-attention module. The latent image after the self-attention module can be calculated as:
$e_{P_{m}}^{i}=A_{r,t}^{R}W_{v}^{i}\mathcal{E}(I_{P_{m}})$ (13)
where $\mathcal{E}$ denotes the image encoder. The output latent of the $i$-th layer in UNet by using region-level adapters is summed as following:
${z^{\prime}}_{t}^{i}=z_{t}^{i}+e_{P_{F}}^{i}+e_{P_{U}}^{i}+e_{P_{L}}^{i},$ (14)
where $e_{P_{F}}$,$e_{P_{U}}$ and $e_{P_{L}}$ denotes the output of each region-level adapter of face region, upper region and lower region, which can be obtained by Eq. 13.
@@
Attention Dynamic Fusion. Mask-based regional-level adapters face the challenge of obtaining region masks that facilitate generation adhering to the layout while remaining faithful to the reference image. To address this, we propose an attention dynamic fusion module to further improve the quality of the generated image. Specifically, we characterize the masks as soft-label masks by removing $M_{r}(x,y)$ mentioned in Eq. 11 and integrate them into region-level adapters to produce coherent image generation. Thus, The output of the region-level adapters can be rewritten as following:
$\begin{split}A_{r,t}^{R}&=\sum\limits_{x=0}^{w}\sum\limits_{y=0}^{h}A_{r,t}(x,y)\\ \end{split}$ (15)
Then $A_{r,t}^{R}$ is passed into Eq. 13-14 to obtain the hidden state of the $i$-th layer in UNet using dynamic region-level adapter.
3.2.3 Multi-character Generation
@@
In addition to achieving remarkable performance in terms of single-character consistency, Character-Adapter also demonstrates remarkable performance for multi-character generation. Similar to single-character generation, we utilize prompt-guided segmentation to automatically localize multiple reference characters. The prompt for $J$ character is noted as:
$P=\{P_{m}^{j}\},m=F,U,L\quad j=1,...,J,$ (16)
@@ 4.1 Experimental setup
Implementation Details. We conduct all experiments on SD v1.5 [30] with classifier-free guidance as 7 and 20-timestep Euler A sampling. The mask threshold $\gamma_{1}$ and $\gamma_{2}$ are all set as 0.8. We use three identity IP-Adapter-Plus in dynamic region-level adapters based on SD1.5. We conducted our experiments using NVIDIA A30 GPUs, more details are provided in Appendix.
@@ 4.2 Quantitative Comparison
Table 1: Quantitative results (%) of Character-Adapter with other methods. Evaluations are conducted for both single and multiple character generation. The best results of training-free methods are highlighted in bold. The best results of finetune-based methods are highlighted in underline.
We conduct experiments on Character-Adapter and previous SOTA works, quantitative results are illustrated in Tab. 1. state-of-the-art performance in zero-shot character consistency generation, with an improvement of 24.8% on the CLIP-I[33] and DINO [34] scores compared to previous state-of-the-art methods in both single and multi character generation. In comparison with fine-tuning-based approaches, Character-Adapter can achieve comparable character consistency while offering over 70 times improvement in computational efficiency, as it requires no additional training. Moreover, Character-Adapter achieves the best performance in terms of text-image alignment with an improvement of 3.5%.
@@ 4.3 Qualitative Comparison
As shown in the multi-character generation example (Fig. 3 6th row), methods like T2I-Adapter, InstantID, Reference-only, and FastComposer fail to preserve the reference characters’ attire. IP-Adapter encounters concept fusion and detail omission issues, attributable to using an image-level adapter, leading to inadequate feature extraction and representation. While fine-tuning-based models like LoRA align with text prompts, they lack details of reference characters and cause concept fusion in multi-character generation. In contrast, Character-Adapter generates high-fidelity multi-character images with intricate details while preserving text-image alignment. Overall, the qualitative results substantiate Character-Adapter’s significant improvements in high-fidelity character generation. Additional results are provided in Appendix.